14 is added to 2/3 of a certain number the results is 5/4 times the original number.find the number

Answer by rfer(16322)

2/3x+14=5/4x
14=5/4x-2/3x
14=15/12x-8/12x
14=7/12x
14/(7/12)=14
14*12/7=x
24=x